OpenStudy (anonymous):

Solve: /2x+5/=7 my answer I got is x=1 and x=-6 is this correct

OpenStudy (anonymous):

@jim_thompson5910 sorry I have another one for you to look at for me

hero (hero):

I'm assuming you mean this: |2x + 5| = 7

hero (hero):

Then 2x + 5 = 7 2x + 5 = -7 2x = -2 2x = -12 x = -1 x = -6 True

OpenStudy (anonymous):

yes @Hero that is what I meant; How did you get -1

hero (hero):

2x + 5 = 7 2x = 7 - 5 2x = 2 x = 1 My bad

OpenStudy (anonymous):

why this is what I got

OpenStudy (anonymous):

oktha

OpenStudy (anonymous):

so it would be true then correct as if you insert -6 back into the problem it does not =7

OpenStudy (anonymous):

not be true I meant

hero (hero):

You have to study what absolute value means in order to understand

OpenStudy (anonymous):

but my solution of -6 and 1 are correct

hero (hero):

Yes, they are correct, now go and study absolute value to find out why.

OpenStudy (anonymous):

thanks